With the following information make a calculation of the direct cost, direct labor and determine the base cost of raw material and the labor base cost according to the following lists.
The company KIA motors SA de CV presents the following materials:
1. #12 steel sheets, $ 1500 each.
2. High quality industrial iron, $ 45,000 per ton
3. Industrial rubber # 23 at a cost of $ 13,500 per ton
4. Electronic sensors package for a unit at a cost of $ 4,500.
5. Set of lights for one unit $ 12,500
6. Engine of 5 horses of 16 valves, 33,500 per unit
7. Game of internal accessories of the vehicle, $ 1,200 per unit package
8. Set of screws for a total assembly, $ 9,800 per game
9. Industrial nylon for protection of the units, $ 4,500 per ton
10. Transfer of finished units to the warehouse, $ 250 per unit

Labour expenses
1. 10 specialized technicians, each one receives a salary of $ 4,500
2.15 auxiliary technicians, each one receives $ 750 per day.
3. A plant supervisor with a monthly salary of $ 35,000
4. 3 auxiliary supervisors with a monthly salary of $ 15,000
5. The technicians generate a social security tax of 26% monthly on salaries.
6. A 5% monthly tax is paid to Health insurance.
7. A monthly payroll tax is paid on the technicians.

8. Finally it takes some unforeseen extra hours for auxiliary technicians of 9 hours per week.

1.Determine the base cost of the car per unit if for the case of the materials 10 cars are produced with the materials of point 2, 3 and 9 and for the case of point 1,4 steel sheets are used, in all the others a set of units.
2. Determine the base cost of labor and consider that for the construction of a car they do it in 3 hours a day.
3. Determine the total production cost if 300 cars are produced per month.
Answer #1

1. Computation of Base cost of car per unit

Sr. No. Particulars Cost ($)    (a) Units    (b) Cost per unit (c = a/b) ($) Units required per car (d) Base Cost of Car per unit (c x d)
MATERIALS
1 Steel Sheets 1500 1 unit 1500 4 6000
2 High Quality Industrial iron 45000 1 ton 45000 1 ton / 10 cars = 0.10 ton 4500
3 Industrial Rubber 13500 1 ton 13500 1 ton / 10 cars = 0.10 ton 1350
4 Electronic sensors package 4500 1 unit 4500 1 unit 4500
5 Set of lights 12500 1 unit 12500 1 unit 12500
6 Engine of 5 horses of 16 valves 33500 1 unit 33500 1 unit 33500
7 Game of internal accessories 1200 1 unit 1200 1 unit 1200
8 Set of screws for total assembly 9800 1 Game 9800 1 Game 9800
9 Industrial Nylon 4500 1 ton 4500 1 ton / 10 cars = 0.10 ton 450
10 Transfer of Finished goods to Warehouse 250 1 unit 250 1 unit 250
TOTAL 74050

2. Base cost of Labor calculations

Given that 300 cars are produced per month. Also that 1 car requires 3 hours of labor.

So, Labor hours per month = 300 cars x 3 hours per car = 900 hours per type of labor.

We'll first compute the labor cost per car for each labor separately.

Specialized Technicians

  • 10 technicians
  • monthly salary = $ 4500 per technician
  • Social Security tax = 26% on Salary = 26% of $4500 = $1170
  • Health Insurance = 5% of Salary = 5% of $4500 = $225
  • Total costs of 1 Specialized technician per month = $4500 + $1170 + $225 = $5895
  • Total costs towards 10 Technicians = $58,950
  • 900 hours of work is done per month.
  • So, Labor cost of Specialized Technicians per hour = $58,950 / 900 hours per month = $65.5 per hour

Auxiliary Technicians

  • In the absence of information of working hours per day and working days per month, we will assume that the auxiliary technicians also work for the same time per day as Specialized technicians work( which is 9 hours per day)
  • So, each auxiliary technician works 9 hours a day
  • 15 technicians (as the pay is per technician per day, number of technicians become irrelevant here, as we will make calculations in terms of number of technician days)
  • Daily pay = $ 750 per technician
  • Work required per month = (3 hours x 300 cars) + (9 unforeseen hours x 4 weeks) = 900 + 36 = 936 hours
  • Number of work days required per month = 936 hours / 9 hours per day = 104 technician days
  • Monthly pay to Auxiliary Technicians = 104 technician days x $750 per day = $78,000
  • Social Security tax = 26% on $78000 = $20,280
  • Health Insurance = 5% of $78,000 = $3900
  • Total costs towards Auxiliary technicians per month = $78000 + $20280 + $3900 = $102180
  • So, Labor cost of Auxiliary Technicians per hour = $102180 / 936 hours per month = $109.17 per hour

Plant Supervisor

  • 1 Supervisor
  • monthly salary = $ 35000 per month
  • Social Security tax = Payable only for technicians, not supervisors
  • Health Insurance = 5% of Salary = 5% of $35000 = $1750
  • Total costs towards 1 Plant Supervisor per month = $35000 + $1750 = $36750
  • 900 hours of work is done per month.
  • So, Labor cost of Plant Supervisor per hour = $36750 / 900 hours per month = $40.83 per hour

Auxiliary Supervisor

  • 3 Supervisors
  • monthly salary = $ 15000 per month
  • Social Security tax = Payable only for technicians, not supervisors
  • Health Insurance = 5% of Salary = 5% of $15000 = $750
  • Total costs towards 3 Auxiliary Supervisors per month = ($15000 + $750) x 3 = $47250
  • 900 hours of work is done per month.
  • So, Labor cost of Plant Supervisor per hour = $47250 / 900 hours per month = $52.50 per hour

Base Cost of Labor per hour

Specialized Technicians $ 65.50 per hour
Auxiliary Technicians $ 109.17 per hour
Plant Supervisor $ 40.83 per hour
Auxiliary Supervisor $ 52.50 per hour

Labor cost per car

Specialized Technicians Auxiliary Technicians Plant Supervisor Auxiliary Supervisor TOTAL Labor cost per Car
Labour cost per hour (a) $ 65.50 per hour $ 109.17 per hour $ 40.83 per hour $ 52.50 per hour
No. of hours per car (b) 3 936/300 = 3.12 3 3
Labor Cost per Car (a x b) $196.50 $340.61 $122.49 $157.50 $817.10

3. TOTAL PRODUCTION COST FOR 300 HOURS

Total Material Cost per Car = $74,050

Total Labor Cost per Car = $817.10

So, Total Production Cost per Car = $74050 + $ 817.10 = $74,867.10

Total production cost for 300 cars = $74,867.10 x 300 = $22,460,130

Please note that the answers will differ if different assumptions are taken, but the calculations proceed in the same method. For example, if health insurance is considered deductible or exempt for social security tax calculations, the answers will differ, ans so on.